Earnings per Share, Price-Earnings Ratio, Dividend Yield The following information was taken from the financial statements ofEarnings per Share, Price-Earnings Ratio, Dividend Yield The following information was taken from the financial statements of Tolbert Inc. for December 31 of the current fiscal year:
Answer #1

a) Earning per share = (668000-300000)/230000 = 1.60 per share

b) Price earning ratio = 13.60/1.60 = 8.50 times

c) Dividend per share = 57500/230000 = 0.25 per share

d) Dividend yield = 0.25*100/13.60 = 1.84%